Order: Ceclor 20 mg/kg/day for an infant who weighs 12 lb 4 oz in divided doses...

  1. Order: Ceclor 20 mg/kg/day for an infant who weighs 12 lb 4 oz in divided doses q8hours.
    1. Give ?
    2. Ceclor is provided as 100 mg / 2 mL.
    3. How many mL would you give?
Answer #1

Ordered Ceclor= 20 mg/kg/day in divided dose q8h

Weight of infant= 12 lb 4 Oz = 12.5 lb=5.68 kg

A) required amount=20X 5.68= 113.6 mg/day in divided doses of q8h

Ceclor available=100mg/2ml

=50 mg/ml

For 113.6 mg required ml= 113.6/50=2.27 ml per day in divided dose of q8h

Per dose ml=2.27/3=0.757 ml